for Quiz #10: Module 5 Problems for oredit s $64.50
Answer #1

1) Net Profit in buying a call option = Share price at expiration - Strike Price - Premium = 54.50 - 51.75 - 2.25 = 0.50

2) Net loss in writing call option =   Strike Price + Premium -Share price at expiration = 51.75 + 2.25 - 54.50 = -0.50

3) Net Loss in buying a put option = Strike Price - Share price at expiration - Premium = 51.75 - 54.75 - 2.25 = -5.25

4) Net Profit in writing a put option =   Share price at expiration + Premium - Strike Price = 54.75 + 2.25 - 51.75  = 5.25

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well